OpenStudy (timaashorty):

What are the Solutions ? 1/2x^2+2x+3 = 0 Please show me step by step, not just the answer Thank you.

OpenStudy (timaashorty):

@Yttrium

OpenStudy (yttrium):

\[\frac{ 1 }{ 2 }x^2 + 2x+3 = 0\] right?

OpenStudy (timaashorty):

right (:

OpenStudy (anonymous):

I will be here to check the answer xD

OpenStudy (timaashorty):

Lol Okay Lola (x

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

what grade is this? :O

OpenStudy (anonymous):

Least common denominator ^_^

OpenStudy (yttrium):

You better get rid of the fraction first. And that is by multiplying the whole equation by the LCD. (Our LCD is 2, know why?)

OpenStudy (timaashorty):

So it could be easier to work out ?

OpenStudy (timaashorty):

@lola honors Algebra 2

OpenStudy (anonymous):

wow :O so timaa, ur older den meh o-o well derp O_O

OpenStudy (yttrium):

Hence. \[2[\frac{1 }{ 2 }x^2+2x+3 = 0]\] \[x^2+4x+6 = 0\] Just factor them out to get the zeroes. And I think this is unfactorable.

OpenStudy (anonymous):

*sits in a corner and thinks* .-.

OpenStudy (yttrium):

yes @timaashorty

OpenStudy (anonymous):

There are none. Your Discriminant, b² - 4ac is -2, so D < 0 If you'd use that in the formula \\[\frac{ -b \pm \sqrt{D}\ }{ 2 }\] it would give: \\[\frac{ -2 \pm \sqrt{-2}\ }{ 2 }\] you'd run into an error because you cannot use the root of a negative factor :)

OpenStudy (anonymous):

Or you can work with getting rid of the 1/2 first to make it easier ;)

OpenStudy (yttrium):

So let's continue. If that is the case, you can use the quadratic formula to get the zeroes. Quadratic formula is: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Where the equation must be in the form of ax^2+bx+c = 0

OpenStudy (yttrium):

Our equation \[x^2+4x+6 = 0\] is already in the form of ax^2+bx+c = 0 Hence, \[\frac{ -4\pm \sqrt{4^2-4(1)(6)} }{ 2(1) }\] Our zeros therefore are: \[\frac{ -4\pm \sqrt{-8} }{ 2 }\] \[\frac{ -4\pm 8i }{ 2 }\] \[-2\pm4i\] --> Final answer.

OpenStudy (timaashorty):

THANK YOU VERRRYY VERRY MUCH! xD @mieel & Yttruim (: I apreciate Your help(:

OpenStudy (timaashorty):

And yes Char, I think I am Older, lol How old are you ? (i4got)